OpenStudy (anonymous):

Find the equation of the line that contains the point (-4, 18) and is parallel to the line that contains (2, -2) and (6, 3) y=mx+b form, do not use decimals. Leave improper fractions improper. help? THis was the original question and this answer, y=(5/4)*x+23

OpenStudy (anonymous):

that answer was wrong

OpenStudy (anonymous):

first fine the slope of the paralel line = 3-(-2) / 6-2 = 5/4 use formula y - y1 = m(x - x1) y - 18 = (5/4) (x - (-4)) y = (5/4) x + 23

OpenStudy (anonymous):

its correct fox

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you guys for your help and patence

OpenStudy (anonymous):

yw